OpenStudy (tainted):

MEDAL WILL BE GIVEN Simplify 3/2 + 2/a^2 - 2/a^3 explain / show work

OpenStudy (anonymous):

is there a right hand side to the equation?

OpenStudy (tainted):

No, that is the full equation

OpenStudy (zehanz):

Let's write it in the equation editor:\[\frac{ 3 }{ 2 }+\frac{ 2 }{ a^2 }-\frac{ 2 }{ a^3 }\]YOu want to write this as one fraction. So you have to choose a new denominator, of which 2, a² and a³ are factors.

OpenStudy (zehanz):

Do you know what number to take for the new denominator?

OpenStudy (tainted):

a^3 as the denominator

OpenStudy (zehanz):

What about the number 2?

OpenStudy (zehanz):

YOu also have to take that into account, so it is 2a³

OpenStudy (tainted):

2 is the variable, so ya

OpenStudy (zehanz):

Now this is what you have:\[\frac{ 3 \cdot ... }{ 2a^3 }+\frac{ 2 \cdot ... }{ 2a^3 }-\frac{ 2 \cdot ... }{ 2a^3 }\]You have to find what goes in place of the dots.

OpenStudy (tainted):

\[\frac{ 3a^3 }{ 2a^3 } + \frac{ 2a }{ 2a^3 } - \frac{ 2 }{ 2a^3 }\]

OpenStudy (zehanz):

I agree with the first one. Second one: there already was a factor 2 in the numerator. The denominator has been multiplied with 2a, so you have to multiply the numerator also with 2a, which makes it 4a. Third one: same thing...